# On Conjecture of Binomial Edge Ideals of Linear Type

### v2

@@ Abstract.
Rees algebra, linear type, Cohen-Macaulay, $d$-sequence, $p$-sequence, binomial edge ideal, tree.
@@ Introduction
is called binomial edge ideal of the graph $G$. We call the binomial $f_{ij}$ associated to an edge $\{i,j\}\in E(G)$ as an edge binomial. In [7], Herzog et al. observed that the binomial edge ideal appears naturally in the study of conditional independence ideals, providing a valuable tool to investigate robustness theory in the context of algebraic statistics. In recent years, researchers have investigated various algebraic properties, including radicality, primality, Gröbner bases, primary decomposition, Cohen-Macaulayness, symbolic powers, among others, of binomial edge ideals in relation to their underlying combinatorial properties; see for partial list [5, 7, 14, 15, 18] and the references therein.
Let $A$ be a standard graded polynomial ring over a field $\mathbb{K}$ and $I\subset A$ be a homogenous ideal. The Rees algebra of an ideal $I\subset A$, is defined as $\mathcal{R}(I)=\{\sum_{i=0}^{n}a_{i}t^{i}\mid n\in\mathbb{N}$ and $a_{i}\in I^{i}\}=\bigoplus_{i\geq 0}I^{i}t^{i}\subset A[t]$, where $I^{0}=A$. $\mathcal{R}(I)$ is a classical algebraic structure that encodes numerous asymptotic properties of that ideal. Indeed, the Rees algebra provides insight into how the powers of an ideal sit inside another Noetherian ring. Moreover, there exists a natural epimorphism from $Sym(I)$ to $\mathcal{R}(I)$, where $Sym(I)$ is the symmetric algebra of $I$. For an ideal $I\subset A$, if $Sym(I)\cong\mathcal{R}(I)$, then $I$ is referred to as an ideal of linear type. The defining equations of the symmetric algebra of an ideal can always be computed from a presentation matrix of the ideal.
@@
Huneke introduced the notion of $d$-sequence in [10]. Huneke, and independently Valla, proved that if the ideal is generated by unconditioned $d$-sequence, then the ideal is of linear type in [9, 20]. Costa [4] established that $d$-sequence condition on $I$ is a sufficient condition for the ideal $I$ to be of linear type. Later, Raghavan [16] provided a simpler proof that $d$-sequence ideals are of linear type. To the best of our knowledge, study of the Rees algebra of binomial edge ideal of linear type are the following. In [11], the authors characterised almost complete intersection binomial edge ideals, a subclass of ideal of linear type. In [1, 2], the authors provided a partial answer for Conjecture A. Particularly, the authors characterized $d$-sequence binomial edge ideals of trees and unicyclic graphs at the level of a sequence of edge binomials. In [13], Kumar characterized binomial edge ideals of linear type for closed graphs.
In this paper, we settle Conjecture A for trees; see Theorem 4.1. To achieve this, we introduce a $p$-sequence in Section 2, as defined in Definition 2.1. In ([3, Corollary 6.2]), Bolognini et al. proved that if $G$ is a bipartite graph, then the binomial edge ideal $J_{G}$ coincides with the parity binomial edge ideal $\mathcal{I}_{G}$ (cf. [12]) and Lovász–Saks–Schrijver ideal $L_{G}(2)$ (cf. [8]), use which we establish the linear type properties of these ideals for trees. In Section 3, we proved that trees and a class of unicyclic graphs form a $p$-sequence, see Theorems 3.1 and 3.2. Finally, in Section 5, we present a counterexample illustrating that the conjecture does not hold for unicyclic graphs in general; refer to Proposition 5.1.
This paper is organized as follows. The following section recalls definitions, notations, and known results required in the rest of the paper. In Section 2, we introduce and study $p$-sequence ideals. In Section 3, we shown that binomial edge ideals of trees and class of unicyclic graphs form a $p$-sequence. We study the Rees algebra of binomial edge ideals of trees in Sections 4. In Section 5, we present a counterexample to Conjecture A for a class of unicyclic graphs.
@@ 1. Preliminaries
  • $e$ is called a pendant edge if one of the vertices is a pendant vertex.
@@
The subsequent results describe the colon ideal operation of an edge binomial on the binomial edge ideal.
@@
[17, Definition 1.1] A relation on an ordered sequence of elements $z_{1},\ldots,z_{n}$ of a commutative ring $A$ is a form of positive degree in the graded polynomial ring $R=A[X_{i}\mid 1\leq i\leq n]$, where elements of $A$ have degree 00 and each $X_{i}$ has degree $1$, that yields 00 when evaluated at $z_{1},\ldots,z_{n}$. Let $\Lambda_{j}$ denote the ideal of $R$ generated by the relations on $z_{1},\ldots,z_{n}$ of degree at most $j$. Let $r$ be the least positive integer such that $\Lambda_{r}=\cup_{j\geq i}\Lambda_{j}$. The invariant $r$ is called the relation type of the ideal $I=(z_{1},\ldots,z_{n})$. An ideal of relation type $1$ is said to be of linear type.
@@
Remark 2.1.
Up to the sequence level, $d$-sequences and $p$-sequences are distinct notions. Consider the following examples:
  1. (1)
@@
To demonstrate that $z_{1},\ldots,z_{n}$ forms a $p$-sequence, it is enough to show that for any $i<i_{1}\leq i_{2}$, $gcd(z_{i},z_{i_{1}}z_{i_{2}})=gcd(z_{i},z_{i_{1}})$. From Equation (2.2) it follows that $gcd(z_{i},z_{i_{1}}^{2})=gcd(z_{i},z_{i_{1}})$ for all $i_{1}>i$. From Equation (2.1) it follows that $gcd(z_{i},z_{i_{1}}z_{i_{2}})\mid gcd(z_{i},z_{i_{1}}^{2})$. Then by Equation (2.2) one has $gcd(z_{i},z_{i_{1}}z_{i_{2}})=gcd(z_{i},z_{i_{1}})$, for all $i<i_{1}<i_{2}$, as desired.
@@
Remark 2.3.
@@ 3.2. Ordering on edges of a unicyclic graph
We claim that an ordered sequence of edge binomials with respect to the edges arranged as described in Subsection 3.1 forms a $p$-sequence. Say $d_{1},\ldots,d_{n}$ be the ordered sequence of edge binomials of $G$ as ordered in Subsection 3.1 and set $d_{0}=0$. Notice that the graph associated with edge binomials $d_{1},\ldots,d_{i}$ intersect with edge binomials associated with $d_{j}$ in at most one vertex, due to the arrangement of the edges of $G$, where $i\in\{0,\ldots,n-1\}$ and $j>i$. Then
@@
for all $10\leq i\leq n-$ 1, and $i<i_{1}\leq i_{2}\leq n$. The above equality is supported by Remark 1.2, since either:
  • The edge binomial associated with $d_{i_{2}}$ and the edge binomial associated with $d_{i_{1}}$ share a common vertex when intersecting with the graph associated with edge binomials $d_{1},\ldots,d_{i}$, or
  • The edge binomial associated with $d_{i_{2}}$ intersects with the graph associated with edge binomials $d_{1},\ldots,d_{i}$ and has a vertex of degree less than or equal to $1$.
Therefore, the edge binomial $d_{i_{2}}$ behaves as a non zero divisor modulo $(d_{0},\ldots,d_{i}):d_{i_{1}}$ as desired. ∎
@@
We claim that an ordered sequence of edge binomials with respect to the edges arranged as in Subsection 3.2 forms a $p$-sequence. Say $d_{1},\ldots,d_{n}$ be the ordered sequence of edge binomials of $G$ and set $d_{0}=0$. Form Theorem 3.1 it follows that $(d_{0},\ldots,d_{i}):d_{i_{1}}d_{i_{2}}=(d_{0},\ldots,d_{i}):d_{i_{1}}$ for all $20\leq i\leq n-$ 2, and $i<i_{1}\leq i_{2}\leq n-1$. It remains to show that
@@
for all $10\leq i\leq n-$ 1, and $i<i_{1}\leq n$. For $i_{1}<n$, the equality follows from Remark 1.2, since edge binomial associated with $d_{n}$, say $e=\{u,v\}$ then $\deg_{H}(u)$ and $\deg_{H}(v)$ is less than or equal to $1$ by the arrangement of edges, where $H$ is the graph associated with edge binomials of the binomial edge ideal $(d_{1},\ldots,d_{i}):d_{i_{1}}$. Note that edge binomial associated with $d_{i_{1}}$ is a bridge in the graph associated with edge binomials $(d_{1},\ldots,d_{i},d_{i_{1}})$. For $i_{1}=n$, the equality follows from Remark 1.1. Hence, $J_{G}$ is a $p$-sequence ideal. ∎
@@ 4. On Conjecture
Let the edge set of $T_{i}$ be $\{e_{i+1},\ldots,e_{n}\}$ for a fixed $i$. Then one has $J_{G_{i_{(e_{i+1})}}}:J_{T_{i}}=J_{G_{i_{(e_{i+1})}}}:f_{e_{i+1}}\cap\cdots\cap J_{G_{i_{(e_{i+1})}}}:f_{e_{n}}$. Note that for any edge $e$ in $T_{i}$, $e$ is a bridge in $G_{i}\cup e$. By the arrangement of edges of $G$, for any edge $e$ in $T_{i}$ which intersect $G_{i}\setminus e_{i}$ has a vertex of degree less than or equal $1$. Then from Remark 1.2 it follows that $J_{G_{i_{(e_{i+1})}}}:f_{e_{j}}=J_{G_{i_{(e_{i+1})}}}$ for all $j=i+1,\ldots,n$. Hence Equation 4.1 holds. The containment $J_{G_{i_{(e_{i+1})}}}:f\supset J_{G_{i_{(e_{i+1})}}}$ follows from Equation 4.1. For the other side containment, suppose $a\notin J_{G_{i_{(e_{i+1})}}}$ such that $af\in J_{G_{i_{(e_{i+1})}}}$, then it follows from Remark 1.3 that there is path from $\alpha$ to $\beta$ in $T_{i}$, where $\{\alpha,\beta\}\in V(G_{i_{(e_{i+1})}})$. Based on the hypothesis that $G$ is a tree, it is not possible, as desired. ∎
@@
Let $f_{e_{1}},\ldots,f_{e_{n}}$ be an ordered sequence of edge binomials for a tree as described in Subsection 3.1. According to Theorem 3.1 the sequence $f_{e_{1}},\ldots,f_{e_{n}}$ form a $p$-sequence. We propose that for any polynomial $F(X_{1},\ldots,X_{n})$ in $R$, if evaluated at $(f_{e_{1}},\ldots,f_{e_{n}})$ results in 00, then $F$ is of linear relation type. The proof is by strong induction on $\deg(F)$. If $\deg(F)=1$, there is nothing to prove. Assume that the statement holds for all $\deg(F)<d$. For $\deg(F)=d$, let $F=\sum_{\lvert\alpha\rvert\leq d}a_{\alpha}X^{\alpha}$ be a polynomial such that evaluated at $f_{e_{1}},\ldots,f_{e_{n}}$ yields 00, where $\alpha\in\mathbb{N}^{n}$. Then, we claim that $F$ is of linear type. Set $F_{0}=F$.
@@
Case I. If $F_{k-1}(X_{1},\ldots,X_{n})=aX_{n-(k-1)}^{\ell_{1}}\cdots X_{n}^{\ell_{k}}$ mod $I_{n-k}$, where $\lvert\ell_{1}\rvert+\cdots+\lvert\ell_{k}\rvert\leq d$. Then it follows 00 mod $\bar{I}_{n-k}$, since $\bar{F}=0$. Thus $af_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}\in\bar{I}_{n-k}$. From Proposition 2.1 and Definition 2.1 it follows that $a\in\bar{I}_{n-k}:f_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}=\bar{I}_{n-k}:f_{e_{n-(k-1)}}$. Thus $af_{e_{n-(k-1)}}\in\bar{I}_{n-k}$. Therefore one has $af_{e_{n-(k-1)}}-\sum_{i=1}^{n-k}b_{i}f_{e_{i}}=0$. Set $H_{k}=aX_{n-(k-1)}-\sum_{i=1}^{n-k}b_{i}X_{i}$ and $F_{2}=F_{1}-H_{2}(X_{n-(k-1)}^{\ell_{1}-1}\cdots X_{n}^{\ell_{k}})$.
Case II. If $F_{k-1}(X_{1},\ldots,X_{n})=aX_{n-(k-1)}H$ mod $I_{n-k}$, where $H\in S[X_{n},\ldots,X_{n-(k-1)}]$, $\deg(H)<d$ and $H$ is not a monomial. Then $\bar{F}_{(k-1)}=af_{e_{n-(k-1)}}\bar{H}$ mod $\bar{I}_{n-k}=0$ mod $\bar{I}_{n-k}$. This implies that $a\bar{H}\in\bar{I}_{n-k}:f_{e_{n-(k-1)}}$. Note that $\bar{I}_{n-k}=J_{G_{n-k}}$ and $e_{n-(k-1)}$ is a bridge in the graph $G_{n-k}\cup e_{n-(k-1)}$. It follows from Remark 1.2 that $a\bar{H}\in J_{G_{n-k_{(e_{n-(k-1)})}}}$. Then from Lemma 4.1 it follows that $a\in J_{G_{n-k_{(e_{n-(k-1)})}}}:\bar{H}=J_{G_{n-k_{(e_{n-(k-1)})}}}$. Therefore, either $a\bar{H}=0$ or every monomial that survive in $a\bar{H}$ belongs to $J_{G_{n-k_{(e_{n-(k-1)})}}}$.
Subcase A. If $a\bar{H}\neq 0$ (i.e, there exists a monomial in the polynomial $aX_{n-(k-1)}H$ which survive on $af_{e}{{}_{n-(k-1)}}\bar{H}$), then choose leading term of $aX_{n-(k-1)}H$ with respect to any term order which survive in $af_{e_{n-(k-1)}}\bar{H}$, say $cX_{n-(k-1)}^{\ell_{1}}\cdots X_{n}^{\ell_{k}}$ mod $I_{n-k}$, where $\lvert\ell_{1}\rvert+\cdots+\lvert\ell_{k}\rvert\leq d$ and $\ell_{1}>0$. Then, we have $cf_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}$ belongs to $\bar{I}_{n-k}$, since $cf_{e_{n-(k-1)}}\in\bar{I}_{n-k}$ as $c\in J_{G_{n-k_{(e_{n-(k-1)})}}}$. Then from Proposition 2.1 and Definition 2.1 it follows that $c\in\bar{I}_{n-2}:f_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}=\bar{I}_{n-2}:f_{e_{n-(k-1)}}$. Therefore one has $cf_{e_{n-(k-1)}}-\sum_{i=1}^{n-k}b_{i}f_{e_{i}}=0$. Set $H_{k,1}=cX_{n-(k-1)}-\sum_{i=1}^{n-k}b_{i}X_{i}$ and $F_{k,1}=F_{k-1}-H_{k,1}(X_{n-(k-1)}^{\ell_{1}-1}\cdots X_{n}^{\ell_{k}})$. Repeat Case II for $F_{2,j}$ until it falls under the case of either Case I or Subcase B for $j\geq 1$.
Subcase B. If $a\bar{H}=0$, then by the induction hypothesis, it follows that there is a linear relation, say $H^{\prime}=\sum_{i=1}^{n}b_{i}X_{i}$ such that $\bar{H^{\prime}}=0$, and $K\in R$ such that $KH^{\prime}=aH$. Then set $F_{k}=F_{k-1}-X_{n-1}KH^{\prime}$.
@@
Now, the claim is reduced to showing that $F_{k}$ is of linear type. By repeating this process, one can conclude that $F$ is of linear type. ∎

### v3

@@ Abstract.
Rees algebra, linear type, $d$-sequence, $p$-sequence, binomial edge ideal, tree.
@@ Introduction
is called binomial edge ideal of the graph $G$. We call the binomial $f_{ij}$ associated to an edge $\{i,j\}\in E(G)$ as an edge binomial. In [7], Herzog et al. observed that the binomial edge ideal appears naturally in the study of conditional independence ideals, providing a valuable tool to investigate robustness theory in the context of algebraic statistics. In recent years, researchers have investigated various algebraic properties, including radicality, primality, Gröbner bases, primary decomposition, among others, of binomial edge ideals in relation to their underlying combinatorial properties; see for partial list [5, 7, 14, 15, 18] and the references therein.
Let $A$ be a standard graded polynomial ring over a field $\mathbb{K}$ and $I\subset A$ be a homogenous ideal. The Rees algebra of an ideal $I\subset A$, is defined as $\mathcal{R}(I)=\{\sum_{i=0}^{n}a_{i}t^{i}\mid n\in\mathbb{N}$ and $a_{i}\in I^{i}\}=\bigoplus_{i\geq 0}I^{i}t^{i}\subset A[t]$, where $I^{0}=A$. $\mathcal{R}(I)$ is a classical algebraic structure that encodes numerous asymptotic properties of that ideal. The symmetric algebra of an ideal $I\subset A$ is defined as $Sym(I)=T(I)/L$, where $T(I)$ is the tensor algebra of $I$ and $L$ is the ideal of $T(I)$ generated by elements of the form $x\otimes_{A}y-y\otimes_{A}x$, where $x,y\in I$. Moreover, there exists a natural epimorphism from Sym$(I)$ to $\mathcal{R}(I)$, where Sym$(I)$ is the symmetric algebra of $I$. For an ideal $I\subset A$, if Sym$(I)\cong\mathcal{R}(I)$, then $I$ is referred to as an ideal of linear type. The defining equations of the symmetric algebra of an ideal can always be computed from a presentation matrix of the ideal.
@@
Huneke introduced the notion of $d$-sequence in [9, 10]. Huneke, and independently Valla, proved that if the ideal is generated by unconditioned $d$-sequence, then the ideal is of linear type in [9, 20]. Costa [4] established that $d$-sequence condition on $I$ is a sufficient condition for the ideal $I$ to be of linear type. Later, Raghavan [16] provided a simpler proof that $d$-sequence ideals are of linear type. To the best of our knowledge, study of the Rees algebra of binomial edge ideal of linear type are the following. In [11], the authors characterised almost complete intersection binomial edge ideals, a subclass of ideal of linear type. In [1, 2], the authors provided a partial answer for Conjecture A. Particularly, the authors characterized $d$-sequence binomial edge ideals of trees and unicyclic graphs at the level of a sequence of edge binomials. In [13], Kumar characterized binomial edge ideals of linear type for closed graphs.
In this paper, we settle Conjecture A for trees; see Theorem 4.1. To achieve this, we introduce a $p$-sequence in Section 2, as defined in Definition 2.1. In Section 3, we proved that trees and a class of unicyclic graphs form a $p$-sequence, see Theorems 3.1 and 3.2. Finally, in Section 5, we present a counterexample illustrating that the conjecture does not hold for unicyclic graphs in general; refer to Proposition 5.1.
@@ 1. Preliminaries
  • An edge $e$ is called a pendant edge if one of the vertices is a pendant vertex.
@@
The subsequent results describe the colon ideal operation of an edge binomial on the binomial edge ideal, which we use repeatedly throughout this article.
@@
[17, Definition 1.1] A relation on an ordered sequence of elements $z_{1},\ldots,z_{n}$ of a commutative ring $A$ is a form of positive degree in the graded polynomial ring $R=A[X_{i}\mid 1\leq i\leq n]$, where elements of $A$ have degree 00 and each $X_{i}$ has degree $1$, that yields 00 when evaluated at $z_{1},\ldots,z_{n}$. Let $\Lambda_{j}$ denote the ideal of $R$ generated by the relations on $z_{1},\ldots,z_{n}$ of degree at most $j$. Let $r$ be the least positive integer such that $\Lambda_{r}=\cup_{j\geq r}\Lambda_{j}$. The invariant $r$ is called the relation type of the ideal $I=(z_{1},\ldots,z_{n})$. An ideal of relation type $1$ is said to be of linear type.
@@
In the subsequent examples, we show that up to the sequence level, $d$-sequences and $p$-sequences are distinct notions.
Example 2.
@@
To demonstrate that $z_{1},\ldots,z_{n}$ forms a $p$-sequence, it is enough to show that for any $i<i_{1}\leq i_{2}$, $\gcd(z_{i},z_{i_{1}}z_{i_{2}})=\gcd(z_{i},z_{i_{1}})$. From Equation (2.2) it follows that $\gcd(z_{i},z_{i_{1}}^{2})=\gcd(z_{i},z_{i_{1}})$ for all $i_{1}>i$. From Equation (2.1) it follows that $\gcd(z_{i},z_{i_{1}}z_{i_{2}})\mid\gcd(z_{i},z_{i_{1}}^{2})$. Then by Equation (2.2) one has $\gcd(z_{i},z_{i_{1}}z_{i_{2}})=\gcd(z_{i},z_{i_{1}})$, for all $i<i_{1}<i_{2}$, as desired.
@@
The following example demonstrates that a $p$-sequence fails to satisfy one of the properties of a $d$-sequence.
Example 4.
@@ 3.2. Ordering on edges of a unicyclic graph
We claim that an ordered sequence of edge binomials with respect to the edges arranged as described in Subsection 3.1 forms a $p$-sequence. Say $d_{1},\ldots,d_{n}$ be the ordered sequence of edge binomials of $G$ as ordered in Subsection 3.1 and set $d_{0}=0$. Let $e_{i}$ be the edge of $G$ associated with the edge binomial $d_{i}$ for all $i$. Now we show that
@@
for all $10\leq i\leq n-$ 1, and $i<i_{1}\leq i_{2}\leq n$.
Let $H$ be the graph formed by the edges $e_{1},\ldots,e_{i}$, for some $i<n$. Notice that the graph $H$ intersect with the edge $e_{j}$ in at most one vertex, due to the arrangement of the edges of $G$, where $i\in\{0,\ldots,n-1\}$ and $j>i$. From Remark 1.2 if follows that $(d_{0},\ldots,d_{i}):d_{i_{1}}=J_{H_{(e_{i_{1}})}}$.
There are three possible way in which the edge $e_{i_{2}}=\{u,v\}$ may intersect with the graph $H$ and the edge $e_{i_{1}}$: Case I. If $e_{i_{2}}\cap H=\emptyset$, then it follows from Remark 1.2 that $J_{H_{(e_{i_{1}})}}:d_{i_{2}}=J_{H_{(e_{i_{1}})}}$. Case II. If $e_{i_{2}}\cap H\neq\emptyset$, and $e_{i_{2}}\cap e_{i_{1}}=\emptyset$, then by the arrangement of the edges of $G$ it follows that $|N_{H_{(e_{i_{1}})}\cup e_{i_{2}}}(u)|=2$ and $|N_{H_{(e_{i_{1}})}\cup e_{i_{2}}}(v)|=1$. Therefore, by applying Remark 1.2 one yields $J_{H_{(e_{i_{1}})}}:d_{i_{2}}=J_{H_{(e_{i_{1}})}}$. Case III. If $e_{i_{2}}\cap H\neq\emptyset$, and $e_{i_{2}}\cap e_{i_{1}}\neq\emptyset$, then either $e_{i_{2}}=e_{i_{1}}$ or share a common vertex, in both the situation neighborhoods of $e_{i_{2}}$ and $e_{i_{1}}$ in $H$ are equal. Thus, the equality follows from Remark 1.2. Hence, the edge binomial $d_{i_{2}}$ behaves as a non zero divisor modulo $(d_{0},\ldots,d_{i}):d_{i_{1}}$ as desired. ∎
@@
We claim that an ordered sequence of edge binomials with respect to the edges arranged as in Subsection 3.2 forms a $p$-sequence. Say $d_{1},\ldots,d_{n}$ be the ordered sequence of edge binomials of $G$ and set $d_{0}=0$. Let $e_{i}$ be the edge of $G$ associated with the edge binomial $d_{i}$ for all $i$. From Theorem 3.1 it follows that $(d_{0},\ldots,d_{i}):d_{i_{1}}d_{i_{2}}=(d_{0},\ldots,d_{i}):d_{i_{1}}$ for all $20\leq i\leq n-$ 2, and $i<i_{1}\leq i_{2}\leq n-1$. It remains to show that
@@
for all $10\leq i\leq n-$ 1, and $i<i_{1}\leq n$.
Observe that if $i<n$, then the edge $e_{i_{1}}=\{r,s\}$ is a leaf in the graph formed by edges $e_{1},\ldots,e_{i},e_{i_{1}}$, and the edge $e_{n}=\{u,v\}$ is a leaf in the graph formed by edges $e_{1},\ldots,e_{i},e_{n}$. Furthermore, we may assume that $s$ and $v$ are pendent vertices in their respective graphs.
For $i_{1}<n$, the equality can be obtained by applying Remark 1.2, since the edge $e_{n}$ is a leaf attached to the graph $H_{(e_{i_{1}})}$, where $H$ denotes the graph formed by the edges $e_{1},\ldots,e_{i}$. We now discuss the possible cases: Case I. If $e_{n}\cap H=\emptyset$, then the equality follows from Remark 1.2. Case II. If $e_{n}\cap H\neq\emptyset$, then there exists an edge $e_{j}$, with $j\leq i$, such that $e_{n}\cap e_{j}\neq\emptyset$. In this case, the vertex $u$ intersect with $e_{j}$. If $deg_{H_{(e_{i_{1})}}\cup e_{n}}(u)=2$ then again the equality follows from Remark 1.2. Suppose if $deg_{H_{(}e_{i_{1})}\cup e_{n}}(u)>2$. Then, it follows from the arrangement of edges of $G$ that $N_{H\cup e_{i_{1}}}(r)=N_{H_{(e_{i_{1})}}\cup e_{n}}(u)$. Now by applying Remark 1.2 one yields the equality.
For $i_{1}=n$, the equality follows from Remark 1.1. Hence, $J_{G}$ is a $p$-sequence ideal. ∎
@@ 4. On Conjecture
Let the edge set of $T_{i}$ be $\{e_{i+1},\ldots,e_{n}\}$ for a fixed $i$. Then one has $J_{G_{i_{(e_{i+1})}}}:J_{T_{i}}=J_{G_{i_{(e_{i+1})}}}:f_{e_{i+1}}\cap\cdots\cap J_{G_{i_{(e_{i+1})}}}:f_{e_{n}}$. Note that for any edge $e$ in $T_{i}$, $e$ is a bridge in the graph $G_{i}\cup e$. For any edge $e=\{u,v\}$ in $T_{i}$ there are three possible ways it may intersect with the graph $G_{i}$ and the edge $e_{i+1}$. Case I. If $e\cap G_{i}=\emptyset$, then it follows from Remark 1.2 that $J_{G_{i_{(e_{i+1})}}}:f_{e}=J_{G_{i_{(e_{i+1})}}}$. Case II. If $e\cap G_{i}\neq\emptyset$ and $e\cap e_{i+1}=\emptyset$, then it follows that $\deg_{G_{i}\cup e}(u)=2$ and $\deg_{G_{i}\cup e}(v)=1$ ($u$,$v$ can be interchangeable). Then, using Remark 1.2 one yields $J_{G_{i_{(e_{i+1})}}}:f_{e}=J_{G_{i_{(e_{i+1})}}}$. Case III. If $e\cap G_{i}\neq\emptyset$ and $e\cap e_{i+1}\neq\emptyset$, then either $e=e_{i+1}$ or $e$ and $e_{i+1}$ shares a common vertex, in both the situation neighborhood of $e$ and $e_{i+1}$ in $G_{i}$ are equal. Therefore, by applying Remark 1.2 one yields $J_{G_{i_{(e_{i+1})}}}:f_{e}=J_{G_{i_{(e_{i+1})}}}$. Hence Equation 4.1 holds. The containment $J_{G_{i_{(e_{i+1})}}}:f\supset J_{G_{i_{(e_{i+1})}}}$ follows from Equation 4.1. For the other side containment, suppose $a\notin J_{G_{i_{(e_{i+1})}}}$ such that $af\in J_{G_{i_{(e_{i+1})}}}$, then it follows from Remark 1.3 that there is path from $\alpha$ to $\beta$ in $T_{i}$, where $\{\alpha,\beta\}\in V(G_{i_{(e_{i+1})}})$. Based on the hypothesis that $G$ is a tree, it is not possible, as desired. ∎
Notation 4.2.
@@
Let $f_{e_{1}},\ldots,f_{e_{n}}$ be an ordered sequence of edge binomials for a tree as described in Subsection 3.1. According to Theorem 3.1 the sequence $f_{e_{1}},\ldots,f_{e_{n}}$ form a $p$-sequence. We propose that for any polynomial $F(X_{1},\ldots,X_{n})$ in $R$, if evaluated at $(f_{e_{1}},\ldots,f_{e_{n}})$ results in 00, then $F$ is of linear relation type. The proof is by strong induction on $\deg(F)$. If $\deg(F)=1$, there is nothing to prove. Assume that the statement holds for all $\deg(F)<d$. For $\deg(F)=d$, let $F=\sum_{\lvert\alpha\rvert\leq d}a_{\alpha}X^{\alpha}$ be a polynomial such that evaluated at $f_{e_{1}},\ldots,f_{e_{n}}$ yields 00, where $\alpha\in\mathbb{N}^{n}$. Then, we claim that $F$ is of linear type.
Set $F_{0}=F$. The proof proceeding in following steps as $k$ runs from $1$ to $n-1$. Here, we define $F_{k}$ recursively in Step $k$.
@@
Case I. If $F_{k-1}(X_{1},\ldots,X_{n})=aX_{n-(k-1)}^{\ell_{1}}\cdots X_{n}^{\ell_{k}}$ mod $I_{n-k}$, where $\lvert\ell_{1}\rvert+\cdots+\lvert\ell_{k}\rvert\leq d$. Then it follows 00 mod $\bar{I}_{n-k}$, since $\bar{F}=0$. Thus $af_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}\in\bar{I}_{n-k}$. From Proposition 2.1 and Definition 2.1 it follows that $a\in\bar{I}_{n-k}:f_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}=\bar{I}_{n-k}:f_{e_{n-(k-1)}}$. Thus $af_{e_{n-(k-1)}}\in\bar{I}_{n-k}$. Therefore one has $af_{e_{n-(k-1)}}-\sum_{i=1}^{n-k}b_{i}f_{e_{i}}=0$. Set $H_{k}=aX_{n-(k-1)}-\sum_{i=1}^{n-k}b_{i}X_{i}$ and $F_{k}=F_{k-1}-H_{k}(X_{n-(k-1)}^{\ell_{1}-1}\cdots X_{n}^{\ell_{k}})$.
Case II. If $F_{k-1}(X_{1},\ldots,X_{n})=aX_{n-(k-1)}H$ mod $I_{n-k}$, where $H\in S[X_{n},\ldots,X_{n-(k-1)}]$, $\deg(H)<d$ and $H$ is not a monomial. Then $\bar{F}_{k-1}=af_{e_{n-(k-1)}}\bar{H}$ mod $\bar{I}_{n-k}=0$ mod $\bar{I}_{n-k}$, since $\bar{F}=0$. This implies that $a\bar{H}\in\bar{I}_{n-k}:f_{e_{n-(k-1)}}$. Note that $\bar{I}_{n-k}=J_{G_{n-k}}$ and $e_{n-(k-1)}$ is a bridge in the graph $G_{n-k}\cup e_{n-(k-1)}$. It follows from Remark 1.2 that $a\bar{H}\in J_{G_{n-k_{(e_{n-(k-1)})}}}$. Then from Lemma 4.1 it follows that $a\in J_{G_{n-k_{(e_{n-(k-1)})}}}:\bar{H}=J_{G_{n-k_{(e_{n-(k-1)})}}}$. Therefore, either $a\bar{H}=0$ or every monomial that survive in $a\bar{H}$ belongs to $J_{G_{n-k_{(e_{n-(k-1)})}}}$.
Subcase A. If $a\bar{H}=0$, then by the induction hypothesis, it follows that there is a linear relation, say $H^{\prime}=\sum_{i=1}^{n}b_{i}X_{i}$ such that $\bar{H^{\prime}}=0$, and $KH^{\prime}=aH$ for some $K\in R$. Then set $F_{k}=F_{k-1}-X_{n-(k-1)}KH^{\prime}$.
Subcase B. If $a\bar{H}\neq 0$ (i.e, there exists a monomial in the polynomial $aX_{n-(k-1)}H$ which survive on $af_{e_{n-(k-1)}}\bar{H}$), then choose leading term of $aX_{n-(k-1)}H$ with respect to any fixed term order which survive in $af_{e_{n-(k-1)}}\bar{H}$, say $cX_{n-(k-1)}^{\ell_{1}}\cdots X_{n}^{\ell_{k}}$, where $\lvert\ell_{1}\rvert+\cdots+\lvert\ell_{k}\rvert\leq d$ and $\ell_{1}>0$. We show that each such surviving monomial can be linearized. Then, the monomial $cf_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}$ belongs to $\bar{I}_{n-k}$, since $cf_{e_{n-(k-1)}}\in\bar{I}_{n-k}$ as $c\in J_{G_{n-k_{(e_{n-(k-1)})}}}$. Then from Proposition 2.1 and Definition 2.1 it follows that $c\in\bar{I}_{n-k}:f_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}=\bar{I}_{n-k}:f_{e_{n-(k-1)}}$. Therefore one has $cf_{e_{n-(k-1)}}-\sum_{i=1}^{n-k}b_{i}f_{e_{i}}=0$. Set $H_{k,1}=cX_{n-(k-1)}-\sum_{i=1}^{n-k}b_{i}X_{i}$ and $F_{k,1}=F_{k-1}-H_{k,1}(X_{n-(k-1)}^{\ell_{1}-1}\cdots X_{n}^{\ell_{k}})$. Repeat Case II for $F_{k,1}$. If it falls under either Case I or Subcase A, then proceed to step $k+1$. Otherwise, recursively define $H_{k,j}$ and $F_{k,j}$, and continue applying Case II to $F_{k,j}$ for $j\geq 2$, until it eventually falls under either Case I or Subcase A. This recursive process terminates in finitely many steps, as $aX_{n-(k-1)}H$ is a polynomial.
@@
Observe that at each kth step, the claim reduces to showing that $F_{k}$ is of linear type. Hence, by the end of Step $n-1$, one can conclude that $F$ is of linear type. ∎
@@
In ([3, Corollary 6.2]), Bolognini et al. proved that if $G$ is a bipartite graph, then the binomial edge ideal $J_{G}$ coincides with the parity binomial edge ideal $\mathcal{I}_{G}$ (cf. [12]) and Lovász–Saks–Schrijver ideal $L_{G}(2)$ (cf. [8]), use which we establish the linear type properties of these ideals for trees in the following remark.